OpenStudy (anonymous):

if the cost of 24lb of yams is 13.32, what will 14 lb cost?

OpenStudy (anonymous):

7.77

OpenStudy (anonymous):

I dont see how you got that answer

OpenStudy (anonymous):

24 lb=13.32 1 lb=13.32/24=.555 14 lb=.555*14=7.77

OpenStudy (anonymous):

it's 7.77, using ratio and proportion. .